Compute for the resultant vector in the given situations:
1. A boy runs 120 m north, then turns 80 m south.
2 An ant moves 5 cm east, then continues to move 14 cm in the same
direction.

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


@ikleyn please don't be rude
If it was obvious to them, they wouldn't have posted. Keep in mind people of all grade levels come on this site.

Problem 1

Think of an xy coordinate grid.
Start at the origin. Then move up 120 units to represent going 120 meters north.

Then move down 80 units to arrive at 120-80 = 40 on the y axis.

The result or final location is that we're 40 meters north of the starting point. For more information, check out the concept about displacement (in physics settings).

Answer: 40 meters north


=====================================================
Problem 2

Once again draw an xy grid, or use one already pre-drawn for you.

Start at the origin. Move 5 units to the right to represent going 5 cm east. Then move another 14 units to arrive at the final location of 5+14 = 19 on the x axis number line.

When writing the vector, don't forget to include information about its direction. This type of problem also deals with displacement.

Answer: 19 cm east

Answer by ikleyn(52873)

.

Hello,   isn't it  OBVIOUS ?


I used to think that for normal person it is as obvious as   2 + 1 = 3.